# Stochastic Approximation with Block Coordinate Optimal Stepsizes

### v1

@@ 5.2.3 Analysis for not using decoupled weight decay
    First, with $\lambda=0$, we have the following inequality in place of (37) (see Lemma 5.1):
@@
    Almost sure convergence as in Theorem 5.4 requires a slightly stronger condition
@@
    To obtain sublinear rates of convergence as in Theorems 5.2 and 5.3, we need an even stronger aiming condition, specifically,
@@
Finally, we comment on the case of adding the regularization $(\lambda/2)\|x\|^{2}$ with $\lambda>0$ but without using decoupled weight-decay. Specifically, we compute $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$ with $\lambda>0$ and apply the BCOS method (48) with $d_{t}$ being $g_{t}$ or its EMA. It is worth emphasizing that, in this case, the aiming condition (50) does not automatically hold with $\mu=\lambda$. This has to do with the coordinate-wise scaling of $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}]}$. To see this, we consider the loss function $F(x)$ defined in (27), i.e.,
$F(x):=\mathbf{E}_{\xi}[f(x,\xi)]+\frac{\lambda}{2}\|x\|^{2},$
and assume that $\mathbf{E}_{\xi}[f(x,\xi)]$ is convex. We contrast the two cases discussed in Section 5.1.1.
  • With full coordinate block, it is straightforward to show that under the convex assumption, the aiming condition (35) is strengthened to
    $\langle x-x_{*},\,\nabla F(x)\rangle\geq\lambda\|x-x_{*}\|^{2}.$
  • However, the above argument does not extend to the case with coordinate-wise scaling. As an example, consider the special case of having uniform SiF across all coordinates. In this case, the corresponding aiming condition (36) cannot be strengthened to
    $\langle x-x_{*},\,\textup{sign}(\nabla F(x))\rangle\geq\lambda\|x-x_{*}\|^{2},$
    because the $\textup{sign}(\cdot)$ function is indifferent or insensitive to the curvature brought by the regularization $(\lambda/2)\|x\|^{2}$. In general, with coordinate-wise scaling and nonunifom SiF, the condition (50) may hold with some positive $\mu$ that is much smaller than $\lambda$.
The above discussion gives a plausible explanation to the empirical observation that decoupled weight decay works better for methods with coordinate-wise scaling (e.g., see Figure 3).
@@ 5.3 Analysis for practical BCOS
Assumption B (Bias of second-moment estimator).
@@
5.3.1 Bounding the difference between conceptual and practical BCOS updates
The following lemma derives an upper bound on the first term on the right-hand side of (52). The proof is given in Appendix B.2.
Lemma 5.5.
Under Assumption B, it holds that:
@@ 5.3.1 Bounding the difference between conceptual and practical BCOS updates
To bound the second term on the right-hand side of (52), we need the following lemma, whose proof is given in Appendix B.3.
Lemma 5.6.
@@
Applying Lemma 5.6 with $Y=d_{t}$ and $Z=v_{t}+\epsilon$, we get the coordinate-wise inequality
@@
With the definition of SNRs, the term containing $\textup{Cov}_{t}(d_{t},v_{t}+\epsilon)$ in (56) can be expressed as
@@
We are ready to bound the difference between the expected updates of the practical BCOS and the conceptual BCOS methods. Specifically, bounding the first term on the right-hand side of (52) using Lemma 5.5 and the second term using the last inequality above, we obtain
@@
$\quad+\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{v_{t}+\epsilon}}\right|\odot\left|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}+\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}+\epsilon)}\right)\right|.$
We can further merge the two terms on the right-hand side by bounding $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{v_{t}+\epsilon}}\right|$ in terms of $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$. To this end, we use Lemma 5.5 to obtain
@@
Lemma 5.7.
Under Assumptions B, we have for all $t\geq 0$,
@@ 5.3.2 Almost sure convergence of practical BCOS
Lemma 5.8.
Under Assumptions B, it holds that
@@
Proof.
Adding and subtracting the term $\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}$ from the inner product, we obtain:
@@
[…]
where the first inequality is due to Hölder’s inequality and the last inequality is due to Lemma 5.7. In addition, we notice that
@@
which concludes the proof. ∎
Theorem 5.4 (Almost sure convergence of BCOSW).
Consider the practical BCOSW method (28). Suppose that Assumptions A and B hold, $d_{t}$ is bounded almost surely, and $\{\alpha_{t}\}$ satisfies (41). Then there exists $\delta>0$ such that we have with probability one,
@@
The proof of Theorem 5.4 is based on a classical result of stochastic approximation due to Dvoretzky DermanSacks1959 . Here we present an extension that best fits our purpose.
Theorem 5.5 (An extension of Dvoretzky’s Theorem).
@@
There are many extensions of Dvoretzky’s original results Dvoretzky1956 . Theorem 5.5 is a minor variation of Venter (Venter1966, , Theorem 1). More concretely,
  • Theorem 1 of Venter Venter1966 has the sequence $\{a_{t}\}$ being a constant sequence, i.e., $a_{t}=a_{\infty}$ for all $t\geq 0$. The extension to a non-constant sequence $\{a_{t}\}$ is outlined in the original work of Dvoretzky Dvoretzky1956 and admits a simple proof due to Derman and Sacks DermanSacks1959 .
  • Theorem 1 of Venter Venter1966 does not include the sequence $\{h_{t}\}$. The extension with $\sum_{t=0}^{\infty}h_{t}<\infty$ is straightforward based on a simple argument of Dvoretzky Dvoretzky1956 .
  • More generally, the sequences $\{a_{t}\}$, $\{b_{t}\}$, $\{c_{t}\}$, $\{h_{t}\}$ can be non-negative measurable functions of $x_{0},\ldots,x_{t}$, and the conclusion of Theorem 5.5 holds if $a_{\infty}$ is an upper bound on $\limsup_{t\to\infty}a_{t}(x_{0},\ldots,x_{t})$ uniformly for all sequences $\{x_{t}\}$ DermanSacks1959 ; ROBBINS1971 .
Proof of Theorem 5.5.
We first write the practical BCOSW algorithm in the form of (62):
@@
Using Lemma 5.8 and then the aiming condition (39), we obtain
@@
Together with $10<1-\alpha_{t}\lambda<$ 1, we conclude that
@@
In order to apply Theorem 5.5, we are left to check the conditions in (64). Using the assumption $\alpha_{t}\to 0$, the definition of $a_{t}$ in (67) implies that $\lim_{t\to\infty}a_{t}=\delta^{2}$. The conditions on $\{b_{t}\}$ and $\{h_{t}\}$ are satisfied due to the assumption $\sum_{t=0}^{\infty}\alpha_{t}^{2}\leq\infty$. For $\{c_{t}\}$, if $\sum_{t=0}^{\infty}c_{t}<\infty$, then we must have $\|x_{t}-x_{\star}\|^{2}\to 0$ almost surely because of the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, and the conclusion of the theorem holds trivially. Otherwise, $\sum_{t=0}^{\infty}c_{t}=\infty$ allows all the conditions in (64) to hold, so we can invoke Theorem 5.5 to finish the proof. ∎
Following the last remark on Dvoretzky’s theorem on page 5.3.2, the radius of the neighborhood $\delta$ can be refined as follows. Let $\delta_{t}$ be the smallest constant such that $\|x_{t}-x_{*}\|\geq\delta_{t}$ implies
$2\|x_{t}-x_{*}\|_{1}\bigl{(}c_{t}+\mathcal{O}(\epsilon)\bigr{)}\leq\lambda\|x_{t}-x_{*}\|^{2},$
@@ 5.4 Bias and variance trade-off
Theorem 5.4 states that the sequence $\{x_{t}\}$ generated by BCOSW converges almost surely to a neighborhood of $x_{*}$, and the radius of the neighborhood is determined by the magnitude of $\sigma_{t}$ and $\epsilon$. Specifically, smaller $\sigma_{t}$ and $\epsilon$ ensure a smaller neighborhood of convergence. According to (57), the magnitude of $\sigma_{t}$ depends on both the bias and variance of the estimator $v_{t}$. In this section, we examine the delicacy of this bias-variance trade-off and illustrate its effects through several examples.
@@
  • Low bias. This means that Assumption B holds with small $\tau$ and $\epsilon$.
@@
    In this case, Assumption B is satisfied with $\tau=0$ and $\epsilon=0$. Then we have
@@
    which corresponds to the sign-gradient method (21) or the sign-momentum method (22), depending on the choice of $d_{t}$. Among them, sign-momentum is preferred because $\textup{Var}_{t}(m_{t}^{2})$ is usually much smaller than $\textup{Var}_{t}(g_{t}^{2})$.
@@
    In this case, we need large $\tau$ and $\epsilon$ for Assumption B to hold. Correspondingly
@@
    It is possible to derive a sensible bound on the bias with a smoothness assumption on the loss function, but we do not delve into it here. On the other hand, we have a simple expression for its variance (derivation given in Appendix B.4):
@@
    Again, deriving a sensible upper bound on the bias can be quite involved and we do not pursue it here. As for the variance, we get (see Appendix B.4)
@@
    The conditional estimator of BCOS-c. The BCOS-c method (Algorithm 3) also uses $m_{t}$ as the search direction. Instead of relying on EMA, it uses a conditional estimator for $\mathbf{E}_{t}[m_{t}^{2}]$, i.e., $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$, where $\beta$ is the smoothing factor for computing $m_{t}$. We have simple expressions for both the bias and the variance:
@@
    Notice that $|m_{t-1}-\mathbf{E}_{t}[g_{t}]|$ is the bias of using $m_{t-1}$ to estimate $\mathbf{E}_{t}[d_{t}]$, and the bias of $v_{t}$ can be much smaller than that. Meanwhile, the variance of BCOS-c is the same as Adam when setting $\beta_{2}$ in Adam as $\beta_{2}=1-(1-\beta)^{2}$.
@@ 6 Conclusion
BCOS is a family of stochastic approximation methods that exploit the flexibility of using different block-coordinate stepsizes. Rather than using sophisticated ideas from optimization, such as preconditioning, it is based on the simple idea of minimizing the distance from the next iterate to an optimal point. While the optimal stepsizes are not computable, we make several simplifications and focus on constructing efficient statistical estimators (for the second moment of the search direction) in determining the stepsizes. In particular, by leveraging a simple conditional estimator, we derive variants of BCOS that achieve competitive performance against AdamW but require less memory and fewer hyperparameters.
Our convergence analysis builds upon several classical results from the stochastic approximation literature. In particular, we extend the classical aiming condition towards a target point to account for coordinate-wise stepsizes and decoupled weight-decay (interpreted as $L_{2}$-regularization). For the conceptual BCOS method, which assumes exact second moment, we establish almost-sure convergence and $\mathcal{O}(1/t)$ rate of convergence. For practical BCOS methods with various second-moment estimators, we establish almost-sure convergence to a neighborhood of the target point, where the radius of the neighborhood is determined by the bias and variance of the second-moment estimator. This framework covers convergence analysis for a broad family of methods, including the sign stochastic gradient/momentum method, RMSProp and Adam(W). In addition, it provides novel insights that can guide the development of new algorithms.

### v2

@@ 5.2.3 Convergence analysis with $\lambda=0$
    First, with $\lambda=0$, we have the following inequality in place of (35) (see Lemma 1):
@@
    Almost sure convergence as in Theorem 13 requires a slightly stronger condition
@@
    To obtain sublinear rates of convergence as in Theorems 5 and 7, we need an even stronger aiming condition, specifically,
@@ 5.3 Analysis for practical BCOS
Assumption 2 (Bias of second-moment estimator)
@@
5.3.1 Bounding the difference between conceptual and practical updates
The following lemma derives an upper bound on the first term on the right-hand side of (50). The proof is given in Appendix A.2.
Lemma 9
Under Assumption 2, it holds that:
@@ 5.3.1 Bounding the difference between conceptual and practical updates
To bound the second term on the right-hand side of (50), we need the following lemma, whose proof is given in Appendix A.3.
Lemma 10
@@
Applying Lemma 10 with $Y=d_{t}$ and $Z=v_{t}+\epsilon$, we get the coordinate-wise inequality
@@
With the definition of SNRs, the covariance term in (54) can be expressed as
@@
We are ready to bound the difference between the expected updates of the practical BCOS and the conceptual BCOS methods. Specifically, bounding the first term on the right-hand side of (50) using Lemma 9 and the second term using the last inequality above, we obtain
@@
$\quad+\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|\odot\left|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}+\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}+\epsilon)}\right)\right|.$
We can further merge the two terms on the right-hand side by bounding $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|$ in terms of $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$. To this end, we use Lemma 9 to obtain
@@
Lemma 11
Under Assumptions 2, we have for all $t\geq 0$,
@@ 5.3.2 Almost sure convergence of practical BCOS
Lemma 12
Under Assumptions 2, it holds that
@@
Proof Adding and subtracting the term $\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}$ from the inner product, we obtain:
@@
$=$~ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle+\left\langle x_{t}-x_{\ast},\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle$
$\geq$~ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle-\left\|x_{t}-x_{\ast}\right\|_{1}\cdot\left\|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\|_{\infty}$
$\geq$~ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle-\left\|x_{t}-x_{\ast}\right\|_{1}\left(\sigma_{t}\left\|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\|_{\infty}+\mathcal{O}(\epsilon)\right),$
where the first inequality is due to Hölder’s inequality and the last inequality is due to Lemma 11. In addition, we notice that
@@
which concludes the proof.
Theorem 13 (Almost sure convergence of practical BCOS)
Consider the practical BCOS method (29). Suppose that Assumptions 1 and 2 hold, $d_{t}$ is bounded almost surely, and $\{\alpha_{t}\}$ satisfies (39). Then there exists $\delta>0$ such that we have with probability one,
@@
The proof of Theorem 13 is based on a classical result of stochastic approximation due to Dvoretzky1956 and its extension by DermanSacks1959. Here we present a version that best fits our purpose.
Theorem 14 (An extension of Dvoretzky’s Theorem)
@@
There are many extensions of the original result by Dvoretzky1956. Theorem 14 is a minor variation of Venter1966. More concretely,
  • Theorem 1 of Venter1966 has the sequence $\{a_{t}\}$ being a constant sequence, i.e., $a_{t}=a_{\infty}$ for all $t\geq 0$. The extension to a non-constant sequence $\{a_{t}\}$ is outlined in the original work of Dvoretzky1956 and admits a simple proof due to DermanSacks1959.
  • Theorem 1 of Venter1966 does not include the sequence $\{h_{t}\}$. The extension with $\sum_{t=0}^{\infty}h_{t}<\infty$ is straightforward based on a simple argument of Dvoretzky1956.
  • More generally, the sequences $\{a_{t}\}$, $\{b_{t}\}$, $\{c_{t}\}$, $\{h_{t}\}$ can be non-negative measurable functions of $x_{0},\ldots,x_{t}$, and the conclusion of Theorem 14 holds if $a_{\infty}$ is an upper bound on $\limsup_{t\to\infty}a_{t}(x_{0},\ldots,x_{t})$ uniformly for all sequences $\{x_{t}\}$ (DermanSacks1959; ROBBINS1971).
Proof [Proof of Theorem 14] We first write the practical BCOS method in the form of (60):
@@
Using Lemma 12 and then the aiming condition (37), we obtain
@@
Together with $0<1-\alpha_{t}\lambda<1$, we conclude that
@@
In order to apply Theorem 14, we are left to check the conditions in (62). Using the assumption $\alpha_{t}\to 0$, the definition of $a_{t}$ in (65) implies that $\lim_{t\to\infty}a_{t}=\delta^{2}$. The conditions on $\{b_{t}\}$ and $\{h_{t}\}$ are satisfied due to the assumption $\sum_{t=0}^{\infty}\alpha_{t}^{2}\leq\infty$. For $\{c_{t}\}$, if $\sum_{t=0}^{\infty}c_{t}<\infty$, then we must have $\|x_{t}-x_{\star}\|^{2}\to 0$ almost surely because of the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, and the conclusion of the theorem holds trivially. Otherwise, $\sum_{t=0}^{\infty}c_{t}=\infty$ allows all the conditions in (62) to hold, so we can invoke Theorem 14 to finish the proof.
Following the third remark after Theorem 14, the radius of the neighborhood $\delta$ can be refined as follows. Let $\delta_{t}$ be the smallest constant such that $\|x_{t}-x_{*}\|\geq\delta_{t}$ implies
$2\|x_{t}-x_{*}\|_{1}\bigl(\sigma_{t}+\mathcal{O}(\epsilon)\bigr)\leq\lambda\|x_{t}-x_{*}\|^{2},$
@@ 5.4 Bias and variance trade-off
Theorem 13 states that the sequence $\{x_{t}\}$ generated by the practical BCOS method converges almost surely to a neighborhood of $x_{*}$, and the radius of the neighborhood is determined by the magnitude of $\sigma_{t}$ and $\epsilon$. Specifically, smaller $\sigma_{t}$ and $\epsilon$ ensure a smaller neighborhood of convergence. According to (55), the magnitude of $\sigma_{t}$ depends on both the bias and variance of the estimator $v_{t}$. In this section, we examine the delicacy of this bias-variance trade-off and illustrate its effects through several examples.
@@
  • Low bias. This means that Assumption 2 holds with small $\tau$ and $\epsilon$.
@@
    In this case, Assumption 2 is satisfied with $\tau=0$ and $\epsilon=0$. Then we have
@@
    which corresponds to the sign-gradient method (21) or the sign-momentum method (22), depending on the choice of $d_{t}$. Between them, sign-momentum is preferred because $\textup{Var}_{t}(m_{t}^{2})$ is usually much smaller than $\textup{Var}_{t}(g_{t}^{2})$.
@@
    In this case, we need large $\tau$ and $\epsilon$ for Assumption 2 to hold. Correspondingly
@@
    It is possible to derive a sensible bound on the bias with a smoothness assumption on the loss function, but we do not delve into it here. On the other hand, we have a simple expression for its variance (derivation given in Appendix A.4):
@@
    Again, deriving a sensible upper bound on the bias can be quite involved and we do not pursue it here. As for the variance, we get (see Appendix A.4)
@@
    The conditional estimator of BCOS-c. The BCOS-c method (Algorithm 3) also uses $m_{t}$ as the search direction. Instead of relying on EMA, it uses a conditional estimator for $\mathbf{E}_{t}[m_{t}^{2}]$, i.e., $v_{t}=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}m_{t}+(1-\beta)^{2}g_{t}^{2}$, where $\beta$ is the smoothing factor for computing $m_{t}$. The bias and variance can be expressed as:
    $\mathrm{Bias_{t}}(v_{t})$ $=2\beta^{2}(1-\beta)\left|m_{t-1}\odot\bigl(m_{t-1}-\mathbf{E}_{t}[g_{t}]\bigr)\right|,$
    $\textup{Var}_{t}(v_{t})$ $=(1-\beta)^{4}\left(4\beta^{2}m_{t-1}^{2}\textup{Var}_{t}(g_{t})+\,\textup{Var}_{t}(g_{t}^{2})+4\beta m_{t-1}(\mathbf{E}_{t}[g_{t}^{3}]-\mathbf{E}_{t}[g_{t}]\mathbf{E}_{t}[g_{t}^{2}])\right).$
    Notice that $|m_{t-1}-\mathbf{E}_{t}[g_{t}]|$ is the bias of using $m_{t-1}$ to estimate $\mathbf{E}_{t}[d_{t}]$, and the bias of $v_{t}$ can be much smaller than that. For the simple alternative $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$, we have simple expressions for both the bias and the variance:
@@
    The variance of the simple alternative is the same as Adam when setting $\beta_{2}$ in Adam as $\beta_{2}=1-(1-\beta)^{2}$.
@@ 6 Conclusion
BCOS is a family of stochastic approximation methods that exploit the flexibility of using different block-coordinate stepsizes. Rather than using sophisticated techniques from optimization such as preconditioning, it is derived from the simple idea of minimizing the distance from the next iterate to a target point. While the optimal stepsizes are not computable, we make several simplifications and focus on constructing efficient statistical estimators (for the second moment of the search direction) in determining the stepsizes. In particular, by leveraging a simple conditional estimator, we derive variants of BCOS that achieve competitive performance against AdamW but require less memory and fewer hyperparameters.
Our convergence analysis builds upon several classical results from the stochastic approximation literature. In particular, we extend the classical aiming condition towards a target point to account for coordinate-wise stepsizes and decoupled weight-decay. For the conceptual BCOS method, which assumes exact second moment, we establish almost-sure convergence and $\mathcal{O}(1/t)$ rate of convergence. For practical BCOS methods with various second-moment estimators, we establish almost-sure convergence to a neighborhood of the target point, where the radius of the neighborhood is determined by the bias and variance of the second-moment estimator. This framework covers convergence analysis for a broad family of methods, including the sign stochastic gradient/momentum method, RMSProp and Adam(W). In addition, it provides novel insights that can guide the development of new algorithms.